# From Architectures to Applications: A Review of Neural Quantum States

### v2

@@ Abstract
The goal of this article is to give an overview of the current state of NQS, i.e. existing NQS architectures, their training and their performance in quantum state simulation and tomography in comparison to conventional methods. Previously published overviews on neural quantum states can be found in Refs. [20, 38, 39, 40, 40, 41] in the more general context of neural network applications in quantum physics, or more specifically on NQS and their optimization in Refs. [42, 43, 44, 45].
@@ 1 A Short Introduction to Neural Quantum States (NQS)
for a certain basis choice |$\ket{\bm{\sigma}}$ that is given by the $d_{l}$ different local configurations, e.g. the spin configurations or Fock space configurations $\bm{\sigma}$. For a system with $N$ sites, $\bm{\sigma}=(\sigma_{1},\sigma_{2},\dots,\sigma_{N})$ consists of e.g. $\sigma_{i}=0,1$ for spin systems ($d_{l}=2$) or $\sigma_{i}=0,1,2,\dots n_{\mathrm{max}}$ for bosonic systems ($d_{l}=n_{\mathrm{max}}+1$).
@@
which is often split into its amplitude $p_{\bm{\theta}}(\bm{\sigma})=|\psi_{\bm{\theta}}(\bm{\sigma})|^{2}$ and its phase $\phi_{\bm{\theta}}(\bm{\sigma})=\mathrm{Im}\left(\mathrm{log}\,\psi_{\bm{\theta}}(\bm{\sigma})\right)$. To feed an input $\bm{\sigma}$ into the network, $\bm{\sigma}$ is typically one-hot encoded, i.e. the $d_{l}$ different local configurations are encoded binary, resulting in a matrix $\bm{\sigma}\in\mathbb{R}^{N\times d_{l}}$ for every configuration $\bm{\sigma}$ of length $N$. Furthermore, the input is often embedded into a space of dimension $d_{h}$, i.e. using a trainable or physically inspired projection the input is projected onto the $d_{h}$ dimensional space that the network is operating on.
@@
The main difficulty of variational approaches is to come up with a good representation $\psi_{\bm{\theta}}(\bm{\sigma})$ of the true wave function coefficients. Here, the great strength of neural networks, namely their expressive power, comes into play: Neural networks with at least one hidden layer, a sufficient number of parameters and an arbitrary non-linear activation function have the ability to represent continuous functions of any – potentially very complicated – form [21, 22, 23, 24]. This makes them promising candidates for a succesful representation $\psi_{\bm{\theta}}$ that is close to the exact wave function $\psi$. In order to obtain this representation $\psi_{\bm{\theta}}$, the network parameters $\bm{\theta}$ are adjusted during the training of the NQS, i.e. starting from some initialization of the neural network parameters $\bm{\theta}_{0}$, the network parameters are adjusted such that $\psi_{\bm{\theta}}$ approximates the true wave function $\psi$ at the end of the training.
@@
In contrast to most machine learning applications, the training can be done in a self-contained way without the use of external data. In general, the specific design choices for the NQS can have a significant impact on its performance, which will be the focus of Sec. 2: Besides the choice of architecture, e.g. the way how the real and imaginary parts of the wave function coefficients $\psi_{\bm{\theta}}$ are modeled. This can be done by splitting $\psi_{\bm{\theta}}$ into amplitude $p_{\bm{\theta}}$ and phase $\phi_{\bm{\theta}}$ parts and using separate networks or separate output nodes / final layers for each part. Another possibility is to use complex network parameters to model the full $\psi_{\bm{\theta}}$ with a single network. Furthermore, NQS can be designed to obey certain symmetries like $U(1)$ symmetries. The performance of the wave function does moreover depend on the optimization and the specific task under consideration, which is discussed in Sec. 3.
@@
respectively, as well as the Monte Carlo average $\langle\cdot\rangle_{\bm{\sigma}}$. For operators involving only a limited number of matrix elements, $O^{\mathrm{loc}}_{\bm{\theta}}$ can be evaluated very efficiently [46], namely local operators or global operators that do not require the calculation of higher order correlations. The computational cost of Eq. (3) results from the generation of samples $\bm{\sigma}$ from $|\psi_{\bm{\theta}}|^{2}$, as well as from the evaluation of the wave function amplitudes for $\bm{\sigma}$ and its connected samples $\bm{\sigma}^{\prime}$. The computational cost of the former strongly depends on $\psi_{\bm{\theta}}$ being normalized or not, since in the latter case samples can not directly be generated from the wave function and more elaborate approaches like Metropolis sampling are needed. Normalized NQS, using so-called autoregressive architectures, are the topic of Sec. 2.6.
NQS have been shown to be capable of representing a broad range of quantum states [26, 28, 31, 30, 32, 27, 33]. To compare their expressivity to more conventional variational approaches like TNs and PEPS, the relationship between them has been studied [47, 48, 26, 49]. Some NQS architectures have been proven to have strictly the same or higher expressive power than practically usable variational tensor networks [26], see Fig. 1. In particular, a range of works have shown that some NQS can encode volume law states without exponential cost [26, 27, 28, 29, 30], although there are volume-law entangled states like the ground state of the Sachdev-Ye-Kitaev model that can not be represented efficiently [50]. Furthermore, Ref. [51] develops a combination of TNs and autoregressive NQS, which improves the capabilities compared to both the original TN and NQS. However, in contrast to TNs which are guaranteed to converge after a sufficiently long optimization, the training of NQS involves a non-convex landscape. Hence, it can be challenging to find the actual ground state, even when the NQS ansatz itself is expressive enough to capture it. Advanced training strategies to overcome this issue are discussed in Sec. 3.1.
Figure 1: Representative power of Matrix product states (MPS) and neural quantum states (NQS). NQS have been proven to have strictly the same or higher expressive power than practically usable variational tensor networks, see e.g. Ref. [26]. This results from the high connectivity between sites of NQS, here exemplary shown for a two-dimensional recurrent neural network (2D-RNN) as considered in Ref. [49] on the top right, compared to MPS, for which information is passed through the system in a 1D manner (see bottom right). Figures adapted from Refs. [38, 26].
@@ 2 NQS Architectures
Neural network quantum states can be implemented using several techniques, including various neural network architectures and different representations of phase and amplitude parts of the wave function. Each architecture comes with its advantages and specialized training strategies, see also Ref. [52]. Additionally, the choice of architecture can also depend on the physical model under investigation. In this section, we discuss commonly used architectures, their application to physical systems in the literature as well as their advantages and downsides compared to other ansätze.
@@ 2.2 Restricted Boltzmann Machines (RBMs)
with the biases in the visible and hidden layers, $b_{j}$ and $c_{i}$, respectively. The corresponding probability for a given input configuration of physical sites $\sigma$ is given by [60]
@@ 2.3 Convolutional Neural Networks (CNNs) and Group CNNs
Convolutional neural networks (CNNs) are used in processing data with a grid-like topology, most commonly two-dimensional data like images. The building blocks of CNNs are shown in Fig. 4: First, convolutional layers employ filters or kernels to scan the input, which can detect local patterns and capture spatial relationships. Basically, each filter in the network uses the same weights for different parts of the input, making CNNs translationally invariant. This approach significantly reduces the number of parameters compared to fully connected networks. Second, pooling layers downsample the spatial dimensions, reducing computational complexity while preserving important features. The final layer of a CNN typically consists of one or more fully connected layers. The convolution of the input data/feature map $a$ with a given filter or kernel $k$ corresponds to
$[a*k](x)=\sum_{y}\sum_{i=1}a_{i}(y)k_{i}(x-y).$ (10)
@@
The result of this operation is a new feature map. $a_{i}(y)$ refers to the value of the $i-th$ channel of the input feature at position $y$. For an RGB image, for example, there are three channels (red, green, blue), and $a_{i}(y)$ is the intensity of one of these colors at pixel $y$. $k_{i}(x-y)$ is the the value of the $i-th$ channel of the kernel at position $x-y$ within the kernel. The kernel is slid over the input image (or feature maps at later layers), and for each position, the kernel values are multiplied with the corresponding input feature values.
@@
In the context of NQS, CNNs are regarded as a viable approach to deal with the properties of square lattice spin systems. The application of CNNs to solve the highly frustrated $J_{1}-J_{2}$ antiferromagnetic Heisenberg model was first introduced in Ref. [91]. In these systems the sign problem remains a significant challenge for quantum Monte Carlo approaches. Ref. [91] demonstrates how CNNs effectively tackle the challenging problem of finding the ground state of such models. Furthermore, Ref. [30] shows that deep CNNs can encode volume-law entangled states efficiently, requiring only $O(\sqrt{N})$ parameters to represent a 2D system with $N$ particles, instead of $O(N)$ for RBMs or $O(N^{2})$ for fully connected networks. In a typical neural network, adding more layers, i.e. making the network deeper, theoretically allows the network to learn more complex features and improve its performance on tasks. However, in practice, when the network gets too deep, one faces problems such as vanishing gradients, where the gradients (which are used to update the weights in the network during training) become very small and make learning very slow or even stop it entirely. This makes deep CNNs hard to train and often leads to poorer performance. In order to address this gradient issue, Ref. [92, 93] attempted to use deep CNNs for a frustrated $J_{1}-J_{2}$ model on a square lattice using different techniques for network training, including a variant of stochastic reconfiguration tailored for large parameter numbers [93]. Besides these ground state calculations, CNNs have also been applied for dynamics simulations [87, 94, 95], see also Sec. 3.3.
@@
CNNs can also be made autoregressive, allowing for sampling directly from the NQS’ amplitudes instead of more involved Markov chain sampling, see Sec. 2.6. This property can be enforced by masking the CNN inputs $\sigma_{i}$,…, $\sigma_{N}$ for the $i$-th output vector to suppress their contribution to the output, as is done in Ref. [34]. The autoregressive CNN has been used for state reconstruction by [97], see also Sec. 3.6.1.

### v3

@@ Abstract
The goal of this article is to give an overview of the current state of NQS, i.e. existing NQS architectures, their training and their performance in quantum state simulation and tomography in comparison to conventional methods. Previously published overviews on neural quantum states can be found in Refs. [20, 38, 39, 40, 40, 41] in the more general context of neural network applications in quantum physics, or more specifically on NQS and their optimization in Refs. [42, 43, 44, 45]. Furthermore, another review on NQS [46] appeared after the publication of the first version of this work.
@@ 1 A Short Introduction to Neural Quantum States (NQS)
for complex $\psi(\bm{\sigma})$ in a certain basis choice |$\ket{\bm{\sigma}}$ that is given by the $d_{l}$ different local configurations, e.g. the spin configurations or Fock space configurations $\bm{\sigma}$. For a system with $N$ sites, $\bm{\sigma}=(\sigma_{1},\sigma_{2},\dots,\sigma_{N})$ consists of e.g. $\sigma_{i}=0,1$ for spin systems ($d_{l}=2$) or $\sigma_{i}=0,1,2,\dots n_{\mathrm{max}}$ for bosonic systems ($d_{l}=n_{\mathrm{max}}+1$).
@@
which is often split into its amplitude $p_{\bm{\theta}}(\bm{\sigma})=|\psi_{\bm{\theta}}(\bm{\sigma})|^{2}$ and its phase $\phi_{\bm{\theta}}(\bm{\sigma})=\mathrm{Im}\left(\mathrm{log}\,\psi_{\bm{\theta}}(\bm{\sigma})\right)$. To feed an input $\bm{\sigma}$ into the network, $\bm{\sigma}$ can be one-hot encoded, i.e. the $d_{l}$ different local configurations are encoded binary, resulting in a matrix $\bm{\sigma}\in\mathbb{R}^{N\times d_{l}}$ for every configuration $\bm{\sigma}$ of length $N$. Note that this is not necessary for spin-$1/2$ systems where the values of the spins are normally mapped to a sequence of ±$\pm 1$ or $0,1$. Furthermore, the input is often embedded into a space of dimension $d_{h}$, i.e. using a trainable or physically inspired projection the input is projected onto the $d_{h}$ dimensional space that the network is operating on.
@@
The main difficulty of variational approaches is to come up with a good representation $\psi_{\bm{\theta}}(\bm{\sigma})$ of the true wave function coefficients. Here, the great strength of neural networks, namely their expressive power, comes into play: Neural networks with at least one hidden layer, a sufficient number of parameters and an arbitrary non-linear activation function have the ability to represent continuous functions of any – potentially very complicated – form [21, 22, 23, 24]. This makes them promising candidates for a successful representation $\psi_{\bm{\theta}}$ that is close to the exact wave function $\psi$. In order to obtain this representation $\psi_{\bm{\theta}}$, the network parameters $\bm{\theta}$ are adjusted during the training of the NQS, i.e. starting from some initialization of the neural network parameters $\bm{\theta}_{0}$, the network parameters are adjusted such that $\psi_{\bm{\theta}}$ approximates the true wave function $\psi$ at the end of the training.
NQS have been first proposed by Carleo and Troyer in 2017 [25]. In this first work, NQS have been applied for the ground state and dynamics simulations of spin systems. Fig. 1a shows two exemplary results from the original work for the antiferromagnetic Heisenberg (AFH) model: In the top figure, the relative ground state errors $\epsilon_{\mathrm{rel}}$ of the NQS energies $E_{\mathrm{NQS}}$ compared to exact energies $E_{\mathrm{exact}}$, i.e. $\epsilon_{\mathrm{rel}}=(E_{\mathrm{NQS}}-E_{\mathrm{exact}})/|E_{\mathrm{exact}}|$, are shown for a $10\times 10$ square lattice AFH. Already in this early work, NQS achieve competitive results to state-of-the art 2D methods like entangled plaquette states (EPS) and PEPS when using sufficiently many parameters (hidden units). Correspondingly, the NQS dynamics simulations of a spin chain with $80$ sites for two quenches of $J_{z}$ shows very good agreement with $t$-DMRG.
Starting from this paper, many works on NQS have appeared in recent years, exploring different NQS architectures, training approaches and their application to various physical systems. Furthermore, many works concern the theoretical representative power of NQS. NQS have been shown to be capable of representing a broad range of quantum states [26, 28, 31, 30, 32, 27, 33]. To compare their expressivity to more conventional variational approaches like TNs and PEPS, the relationship between them has been studied [47, 48, 26, 49]. Some NQS architectures have been proven to have strictly the same or higher expressive power than practically usable variational tensor networks [26], see Fig. 1. In particular, a range of works have shown that some NQS can encode some volume law states without exponential cost [26, 27, 28, 29, 30], although there are volume-law entangled states like the ground state of the Sachdev-Ye-Kitaev model that can not be represented efficiently [50]. Furthermore, Ref. [51] develops a combination of TNs and autoregressive NQS, which improves the capabilities compared to both the original TN and NQS. However, in contrast to TNs which are guaranteed to converge after a sufficiently long optimization, the training of NQS involves a non-convex landscape. Hence, it can be challenging to find the actual ground state, even when the NQS ansatz itself is expressive enough to capture it. Advanced training strategies to overcome this issue are discussed in Sec. 3.1.
Figure 1: Neural quantum states (NQS): a) In the seminal work bei Carleo and Troyer [25], NQS are used for ground state representations (top) and dynamics simulations of spin systems (bottom), here for antiferromagnetic Heisenberg (AFH) models on a $10\times 10$ lattice (top) and on an $80$ site chain. Figure adapted from Ref. [25]. b) Representative power of Matrix product states (MPS) and NQS. NQS have been proven to have strictly the same or higher expressive power than practically usable variational tensor networks, see e.g. Ref. [26]. This results from the high connectivity between sites of NQS, here exemplary shown for a two-dimensional recurrent neural network (2D-RNN, see blue inset) as considered in Ref. [49] on the top right, compared to MPS, for which information is passed through the system in a 1D manner (see orange inset). Figures adapted from Refs. [38, 26].
@@
In contrast to most machine learning applications, the training can be done in a self-contained way without the use of external data. In general, the specific design choices for the NQS can have a significant impact on its performance, which will be the focus of Sec. 2: Besides the choice of architecture, e.g. the way how the real and imaginary parts of the wave function coefficients $\psi_{\bm{\theta}}$ are modeled. This can be done by splitting $\psi_{\bm{\theta}}$ into amplitude $p_{\bm{\theta}}$ and phase $\phi_{\bm{\theta}}$ parts and using separate networks or separate output nodes / final layers for each part. Another possibility is to use complex network parameters to model the full $\psi_{\bm{\theta}}$ with a single network. The performance of the wave function does moreover depend on the optimization and the specific task under consideration, which is discussed in Sec. 3.
@@
respectively, as well as the Monte Carlo average $\langle\cdot\rangle_{\bm{\sigma}}$. For operators involving only a limited number of matrix elements, namely local operators or global operators that do not require the calculation of higher order correlations, $O^{\mathrm{loc}}_{\bm{\theta}}$ can be evaluated very efficiently [52]. The computational cost of Eq. (3) results from the generation of samples $\bm{\sigma}$ from $|\psi_{\bm{\theta}}|^{2}$, as well as from the evaluation of the wave function amplitudes for $\bm{\sigma}$ and its connected samples $\bm{\sigma}^{\prime}$. The computational cost of the former strongly depends on $\psi_{\bm{\theta}}$ being normalized or not, since in the latter case samples can not directly be generated from the wave function and more elaborate approaches like Metropolis sampling are needed. Normalized NQS, using so-called autoregressive architectures, are the topic of Sec. 2.5. Autoregressive NQS can be designed to obey certain symmetries like $U(1)$ symmetries as discussed in Sec. 2.5.1. For non-autoregressive NQS symmetries can be taken into account in the Monte Carlo sampling.
Exemplary Works and Considered Physical Systems
Architecture Features Ground and Excited States,
Finite $T$ States Open Systems and Dynamics
Feed Forward (frustrated) spin systems [53, 54], spin systems [55]
Neural Networks simplicity bosons [53, 54, 56, 57, 58]
(Sec. 2.1)
(frustrated) spin systems [59, 60, 61, 62, 63, 64],
Restricted well studied, spin liquids [48, 65], topologically spin systems in 1D
Boltzmann e.g. in terms of ordered states [66, 67, 68, 31, 69], [25, 70, 55], ladders [71]
Machines expressivity; bosons [72, 53, 73, 74], and 2D [75, 76, 77]
(Sec. 2.2) interpretability fermions [78], molecules [79]
Convolutional incorporate frustrated spin systems 2D spin systems
Neural Networks lattice [25, 80, 81, 82, 83, 84, 85, 86, 45, 87] on various [75, 88, 89]
(Sec. 2.3) symmetries lattice geometries [90, 91, 92, 93, 94]
Graph applicable to (frustrated) spin systems [95]
Neural Networks any lattice and bosons [96] on various
(Sec. 2.4) geometry lattice geometries
Transformer self-attention (frustrated) spin systems [97, 98, 99, 100], spin systems
Neural Networks mechanism Rydberg states [101, 102, 103], [104, 105, 40]
(Sec. 2.6) quantum chemistry [106, 107, 108, 109]
(frustrated) spin systems
Autoregressive perfect [34, 110, 98, 111, 112, 113], spin glass [114], spin systems
Neural Networks sampling topologically ordered (bosonic) [115, 105],
(Sec. 2.5) states [116, 117], Rydberg states Rydberg states
[118, 101, 103], fermions [119, 35] [120]
Table 1: Overview of NQS architectures and their applications to ground, excited states, dynamics, finite temperature states, and open systems.
Figure 2: Architectures used in the works cited in this review in a time interval of 6 months from 2016 to June 2024: The fraction of works with implementations of variants of feed forward neural networks (FFNNs, Sec. 2.1), restricted Boltzmann machines (RBMs, Sec. 2.2), convolutional neural networks (CNNs, Sec. 2.3), transformers (Sec. 2.6), autoregressive networks (Sec. 2.5), fermionic networks (Sec. 2.7) or other architectures. The total number of works is shown in black; note that for 2024 we consider only works until June 2024. In some works, several architectures are used, and are taken into account as separate works for each architecture here. Autoregressive transformers are counted twice.
@@ 2 NQS Architectures
Neural network quantum states can be implemented using several techniques, including various neural network architectures and different representations of phase and amplitude parts of the wave function. Each architecture comes with its advantages and specialized training strategies, see also Ref. [121]. Additionally, the choice of architecture can also depend on the physical model under investigation.
In this section, we discuss commonly used architectures, their application to physical systems in the literature as well as their advantages and downsides compared to other ansätze. In Fig. 2 the type of architectures used in the works cited in this review are shown. It can be seen that the field started with representations in terms of restricted Boltzmann machines (RBMs, Sec. 2.2), as in the seminal work by Carleo and Troyer in 2017 [25]. In the first two years after that, mainly works with NQS based on convolutional neural networks (CNNs, Sec. 2.3) and feed forward neural networks (FFNNs, Sec. 2.1) appeared, while in recent years autoregressive networks (Sec. 2.5) and transformer neural networks (Sec. 2.6) have gained attention. Furthermore, after a focus on spin systems in the early stage of development of NQS architectures, the field turns towards the simulation of fermionic systems (Sec. 2.7). At the time of publication of this (revised) work (June 2024), works on many different architectures appear with a similar fraction. To the best of our knowledge, this is mostly due to the fact that it is not clear at first sight and which NQS architecture is most suitable for a given physical problem is a major open question in the field. Nevertheless, most architectures have certain strengths, which we attempt to summarize in Tab. 1 and in the remaining part of this section.
@@ 2.2 Restricted Boltzmann Machines (RBMs)
with the biases in the visible and hidden layers, $b_{j}$ and $c_{i}$, respectively. Using Eq. (7) and summing over the hidden nodes $\mathbf{h}$, the corresponding probability for a given input configuration of physical sites $\sigma$ is given by [124]
@@ 2.3 Convolutional Neural Networks (CNNs) and Group CNNs
Convolutional neural networks (CNNs) are used in processing data with a grid-like topology, most commonly two-dimensional data like images. The building blocks of CNNs are shown in Fig. 5: First, convolutional layers employ filters or kernels to scan the input, which can detect local patterns and capture spatial relationships. Basically, each filter in the network uses the same weights for different parts of the input, making CNNs translationally invariant. This approach significantly reduces the number of parameters compared to fully connected networks. Second, pooling layers downsample the spatial dimensions, reducing computational complexity while preserving important features. The final layer of a CNN typically consists of one or more fully connected layers.
Each convolution layer consists of feature maps $a$ and kernels $k$. The kernel $k$ is slid over the input image (or feature maps at later layers), and for each translation $(x-\tilde{x},y-\tilde{y})$, the the kernel values $k_{i}(x-\tilde{x},y-\tilde{y})$ are multiplied with the translated input feature values $a_{i}(\tilde{x},\tilde{y})$. In total, the convolution of the two-dimensional input data/feature map $a$ is
$\tilde{a}(x,y)=[a*k](x,y)=\sum_{\tilde{x},\tilde{y}}\sum_{i}a_{i}(\tilde{x},\tilde{y})k_{i}(x-\tilde{x},y-\tilde{y}).$ (10)
@@
The result of this operation is a new feature map $\tilde{a}$, which is the input of the subsequent layers. The index $i$ refers to the value of the $i-th$ channel of the input feature. For an RGB image, for example, there are three channels (red, green, blue), and $a_{i}(\tilde{x},\tilde{y})$ is the intensity of one of these colors at pixel position $(\tilde{x},\tilde{y})$. $k_{i}(x,y)$ is the the value of the $i-th$ channel of the kernel at position $(x,y)$ within the kernel.
@@
In the context of NQS, CNNs are regarded as a viable approach to deal with the properties of two-dimensional systems. The application of CNNs to solve the highly frustrated $J_{1}-J_{2}$ antiferromagnetic Heisenberg model was first introduced in Ref. [80]. In these systems the sign problem remains a significant challenge for quantum Monte Carlo approaches. Ref. [80] demonstrates how CNNs effectively tackle the challenging problem of finding the ground state of such models. Furthermore, Ref. [30] shows that deep CNNs can encode volume-law entangled states efficiently, requiring only $O(\sqrt{N})$ parameters to represent a 2D system with $N$ particles, instead of $O(N)$ for RBMs or $O(N^{2})$ for fully connected networks. In a typical neural network, adding more layers, i.e. making the network deeper, theoretically allows the network to learn more complex features and improve its performance on tasks. However, in practice, when the network gets too deep, one faces problems such as vanishing gradients, where the gradients (which are used to update the weights in the network during training) become very small and make learning very slow or even stop it entirely. This problem is addressed e.g. in Ref. [87] using skip connections that allow to bypass some layers and layer normalization, where inputs for all neurons within the same layer are normalized for each sample. Furthermore, a variant of stochastic reconfiguration tailored for large parameter numbers is used in this work. Besides these ground state calculations, CNNs have also been applied for dynamics simulations [75, 88, 89], see also Sec. 3.3.